OpenStudy (anonymous):

Find the coordinates of the point (x, y, z) on the plane z = 3 x + 4 y + 4 which is closest to the origin.

OpenStudy (eyust707):

okay so 3-d is very similar to 2-d

OpenStudy (eyust707):



OpenStudy (eyust707):

what point on that line will be closest to the origin?

OpenStudy (eyust707):

you can label it using the edit drawing button if you would like

OpenStudy (accessdenied):

What is the formula for distance from the origin to just some arbitrary point, (x,y,z)? We would use that formula with the constraint that "z = 3 x + 4 y + 4"... from there, I'm not exactly sure how to the answer from there. I guess you'd use lagrange multipliers but I haven't learned about them yet. :P

OpenStudy (dumbcow):

my guess would find normal vector, the closest point will be the one that forms a perpendicular line with the origin. normal = <-3,-4,1> this represents directional vector through origin line represented by: x = -3t y = -4t z = t find point where line intersects plane z = 3x+4y +4 t = -9t -16t +4 26t = 4 t = 2/13 point --> (-6/13, -8/13, 2/13)

OpenStudy (anonymous):

Dumbcow, how did you get t=-9t -16t +4?

OpenStudy (anonymous):

 im guessing around there?

OpenStudy (dumbcow):

i substituted in the parametric form of the line into the plane equation to find the point on the line that is also on the plane

OpenStudy (anonymous):

okay, I see. Thank you!